The demand for beer in Japan is given by the following equation: Qd = 700 − 2P PN + 0.1I, where P is the price of beer, PN is the price of nuts, and I is average consumer income.

a) What happens to the demand for beer when the price of nuts goes up? Are beer and nuts substitutes or complements? - prove

b) What happens to the demand for beer when average consumer’s income rises?

c) Graph the demand curve for beer when PN = 100 and I = 10,000.


1
Expert's answer
2021-05-12T10:17:18-0400

a) The sign in front of the prince of nuts, Pn, is negative. This means when the price of nuts goes up, the beer quantity demanded falls for all levels of price (demand shifts left). Beer and nuts are demand complements.

b) The sign in front of income, I, is positive. This means when income rises, quantity demanded increases for all levels of price (demand shifts rightward).

c) Now:

"Q_d = 700 -2З -100 + 0.1 \times 10000 = 1600 -2P \\

P = 800 -0.5Q_d"

When "Q_d = 0 \; P = 800"

When "P=0 \; Q_d = 1600"